OpenStudy (anonymous):

309600 seconds = ______?_______ hours. A. 86 B. 5,160 C. 18,576,000

OpenStudy (callisto):

309600 seconds = (309600 / 60) minute = (309600 / 60) / 60 hours =??

OpenStudy (anonymous):

Its either A, B or C.

OpenStudy (nali):

it is actually 86

OpenStudy (nali):

becuz 309600 seconds / 60 = 5160/60 = 86

OpenStudy (nali):

do u get it?

OpenStudy (anonymous):

yes thank you.

OpenStudy (nali):

u r welcome